Calculate the maximum and minimum values of resistance given by 900 k ohm +_ 2.5 %
1
Expert's answer
2020-03-09T11:16:25-0400

First we will calculate 2.5% which is equal to "\frac{2.5}{100}*900k" ohm =22.5 k ohm

Maximum Value of resistance = (900k + 22.5k) ohm =922.5k ohm

Minimum Value of resistance =(900k - 22.5k) ohm =877.5k ohm